OpenStudy (anonymous):

will medal and fan!! the endpoints of a line segment are at the coordinates (-6 3 4) and (4 -1 2) what is the midpoint of the segments?

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@Mythorius

OpenStudy (anonymous):

no its on a 3 dimensional plane,

OpenStudy (anonymous):

its fine thanks for trying

OpenStudy (anonymous):

@kitty626oldrose

OpenStudy (anonymous):

@raffle_snaffle

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@mathstudent55

TheSmartOne (thesmartone):

\(\Large\bf \left( \frac{x_1 + x_2 + x_3}{3}, \frac{y_1 + y_2 + y_3}{3}, \frac{z_1 + z_2 + z_3}{3}\right)\)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

never mind that was the wrong formula @TheSmartOne, someone just explained it to me.

TheSmartOne (thesmartone):

Oh my bad. >.< I wasn't thinking. Here is the correct formula: \(\Large\bf \left( \frac{x_1 + x_2}{2}, \frac{y_1 + y_2 }{2}, \frac{z_1 + z_2}{2}\right)\)